OpenStudy (anonymous):

NEED HELP RIGHT AWAY PLEASE HELP WILL FAN Donna's company reimburses her expenses on food, lodging, and conveyance during business trips. The company pays $50 a day for food and lodging and $0.60 for each mile traveled. Donna drove 200 miles and was reimbursed $1,620. Part A: Create an equation that will determine the number of days x on the trip. (3 points) Part B: Solve this equation justifying each step with an algebraic property of equality. (6 points) Part C: How many days did Donna spend on this trip? (1 point)

OpenStudy (jacobciezki):

there saying c

OpenStudy (texaschic101):

50x + 0.60m = 1620 m = 200 50x + 0.60(200) = 1620 50x + 120 = 1620 50x = 1620 - 120 50x = 1500 x = 1500/50 x = 30 check.. 50(30) + 0.60(200) = 1620 1500 + 120 = 1620 1620 = 1620 (correct) so it is 30 days that Donna spent on the trip